OpenStudy (anonymous):

Please help. Im stuck on this bad. Refer to the campground problem from Task 1 on p. 340 of Geometry. Recall that the campsites have the following coordinates: Brighton Bluff at B (2, 2), Ponaganset Peak at P (4, 10), and Harmony Hill at H (12, 2). What are the coordinates of the point where the showers should be placed in order to be equidistant from all three campsites?

OpenStudy (anonymous):

@perl @AriPotta @adilalvi @AriPotta @Abhisar @Blu-Girl @BunnyBree @babygurl47672 @briensmarandache @bkonch @Cman456 @ciara9271 @Catlover5925 @CausticSyndicalist @DivergentMockingjay19 @dudezack @eliassaab @eighthourlunch Can anyone help

OpenStudy (anonymous):

let me see

OpenStudy (anonymous):

I never grasped the orthocener, circumcenter, and that fun stuff.

OpenStudy (anonymous):

I am really sorry but I can't help

OpenStudy (anonymous):

@Ria23 @triciaal @tualatin_wolves01 @tayromo @tualatin_wolves01 @kmac416 @karimse07 @kitykat9 @KaiFifer @Write0The0Right0For0Wrong0Doing @walters @linn99123 @Librarian @lexycollins Anyone know this?

OpenStudy (catlover5925):

i sent you a message @Beleaguer

OpenStudy (anonymous):

@ganeshie8 @goformit100 @gleem @Godlovesme @One098 @SolomonZelman @sourwing

OpenStudy (briensmarandache):

i think this is the logic to follow

OpenStudy (catlover5925):

here ya go this is how i found it

OpenStudy (anonymous):

Theres an equation that goes along with it. It has something to do with finding the slop of two points and I never understood it.

OpenStudy (catlover5925):

ok what is this assignment called i will try to google an equation for it

OpenStudy (anonymous):

Inequalities of a triangle.. maybe. Try looking for sonething that has to do with concurrency of a triangle with points. Thats what im looking for

OpenStudy (catlover5925):

is this the equation??

OpenStudy (anonymous):

Nope. At least im not the only one with trouble in this problem haha.

OpenStudy (briensmarandache):

have you tried my solution?

OpenStudy (anonymous):

Yea, but did you use an equation other than just plotting the points and finding the middle? That is what its asking me.

OpenStudy (briensmarandache):

this is what i got, just needs to be simplified

OpenStudy (catlover5925):

@briensmarandache did it make sense what i did??

OpenStudy (briensmarandache):

i wasnt sure what you were showing, i just expressed what i told her to do.

OpenStudy (anonymous):

To check your answer, the shower should be at (7,5), the blue point in the attached graph

OpenStudy (anonymous):

You really have to find the center of the circle passing by the three given points.

OpenStudy (briensmarandache):

no, if you look at the attachment you will see mid point BHPH, that is the center

OpenStudy (briensmarandache):

i have to leave for work if your still stuck ill be on around ten

OpenStudy (anonymous):

Yes, if the points are {{2, 2}, {4, 10}, {12, 2}} then my answer is correct. See the picture

OpenStudy (anonymous):

@briensmarandache , did you look at the attached circle?

OpenStudy (briensmarandache):

well my answer is different so im not sure by my logic i feel as if im correct but i could be wrong

OpenStudy (anonymous):

Let us find the center of the circle that goes to the points {{2,2},{4,10},{12,2}}. Let \[ (x-a)^2 + (y-b)^2 =c\] where c is the square of the radius We need 3 equations let us write them down \[ (2-a)^2+(2-b)^2 =c\\ (4-a)^2+(10-b)^2=c\\ (12-a)^2+(2-b)^2=c \] Wolframalpha will solve them for us to get http://www.wolframalpha.com/input/?i=solve%28+%282-a%29%5E2%2B%282-b%29%5E2+%3Dc%2C+%284-a%29%5E2%2B%2810-b%29%5E2%3Dc%2C+%2812-a%29%5E2%2B%282-b%29%5E2%3D+c+%29 a=7,b=5, c=34

OpenStudy (anonymous):

Actually, you can see geometrically why a =7

OpenStudy (anonymous):

Thanks for the answer! Your aweomse :) *throws medal*